2. It is known that, when given a choice between ordering an espresso or a cappuccino, 63% of people will order the cappuccino. For a sample of 253 people, what is the probability that sampling error for the proportion will be greater than 0.05?


1
Expert's answer
2021-11-01T07:07:22-0400

sampling error:


"z\sqrt{\frac{p(1-p)}{n}}>0.05"


"z\sqrt{\frac{0.63(1-0.63)}{253}}=0.03z>0.05"


"z>5\/3=1.67"


the probability that sampling error for the proportion will be greater than 0.05:


"P(z>1.67)=1-P(z<1.67)=1-0.9525=0.0475"